### v1

# DocFinQA: A Long-Context Financial Reasoning Dataset

Abstract
Research in quantitative reasoning within the financial domain indeed necessitates the use of realistic tasks and data, primarily because of the significant impact of decisions made in business and finance. Financial professionals often interact with documents hundreds of pages long, but most research datasets drastically reduce this context length. To address this, we introduce a long-document financial QA task. We augment 7,621 questions from the existing FinQA dataset with full-document context, extending the average context length for each question from under 700 words in FinQA to 123k words in DocFinQA. We conduct extensive experiments of retrieval-based QA pipelines and long-context language models on the augmented data. Our results show that DocFinQA provides challenges for even the strongest, state-of-the-art systems.
1 Introduction
Automated financial analysis is challenging because it involves reasoning over large volumes of textual and tabular data Azzi et al. (2019). Existing works on automating financial numerical reasoning focus on unrealistically specific document snippets Chen et al. (2021); Zhu et al. (2021). Datasets are often limited to narrowly focused document sections, failing to reflect the broader and more realistic scenarios faced by analysts Masson and Montariol (2020). Financial professionals typically sift through hundreds of pages per document, requiring deep understanding of both the content and structure to effectively navigate and extract pertinent information. Current long-document QA datasets such as Kočiský et al. (2018) do not test the quantitative reasoning skills needed in the financial domain.
In this work, we introduce DocFinQA, a new long-document financial question answering task. We extend the FinQA dataset of expert annotated questions and answers Chen et al. (2021) with full Securities and Exchange Commission (SEC) reports. This results in a significantly larger context in the DocFinQA dataset – by an average factor of 175 – than the context in the FinQA dataset. The resulting long-document QA task offers a more robust evaluation of a model’s reasoning capabilities over long contexts. In line with recent work on program synthesis for financial QA Koncel-Kedziorski et al. (2023), the questions in DocFinQA are annotated with Python programs to generate the answers, allowing for training and evaluating program synthesis models for use in realistic financial workflows.
Figure 1: DocFinQA extends FinQA to inputs with over 100K tokens. The resulting dataset is more challenging, causing GPT-4 to make nearly twice as many errors.
Using this new setup, we conduct detailed experiments of retrieval-based and long-context LLM systems. We study a typical retrieval pipeline consisting of chunking and encoding the document, searching for the best chunks given a question, and passing a pair of a question and top-$k$ chunks to a generative QA model Hsu et al. (2021). We also evaluate retrieval-free approaches using long-context LLMs Weston and Sukhbaatar (2023). Our results show that for LLMs to be effective in financial settings it is necessary to directly address the specific nuances of the financial domain, like large contexts; our dataset effectively tests these nuances.
2 Related Work
Prior studies in financial question answering focus on non-numerical reasoning Day and Lee (2016); Jørgensen et al. (2023); Maia et al. (2018). Short-context grounded numerical reasoning tasks were introduced with datasets such as FinQA Chen et al. (2021) and TAT-QA Zhu et al. (2021). Recently, understanding long documents has attracted more attention for tasks involving events Yang et al. (2018), table of contents Bentabet et al. (2020), and causal relations Mariko et al. (2022). However, to the best of our knowledge, ours is the first to address financial numerical QA by grounding it with extremely long documents – upwards of hundreds of pages of context for a given question.
Long document QA has been studied in NLP with the introduction of datasets such as SearchQA Dunn et al. (2017), NarrativeQA Kočiský et al. (2018), and QuALITY Pang et al. (2022). Due to the limited context size of LLMs, retrieval-based models are commonly used to filter irrelevant text Izacard et al. (2022); Lewis et al. (2020). Recently, advances in attention mechanisms Beltagy et al. (2020); Dao et al. (2022) and positional embeddings Press et al. (2021); Su et al. (2023) allow for conducting end-to-end grounded QA with context windows of more than 100k tokens. However, these methods suffer from loss of important context Zhang et al. (2023) and often fail to make full use of longer inputs Liu et al. (2023). Our work studies the intersection of numerical reasoning and long document processing, and our results show that there is still much room for improvement.
Dataset #Docs #QAs Multi-page Number Table #Words
[…]
Table 1: Comparison of DocFinQA and existing datasets for Finance QA and Long Document QA.
3 DocFinQA Dataset
To assess the reasoning capabilities of LLMs for long-context financial question answering, we take the FinQA dataset as our starting point. FinQA consists of questions created by financial experts. These questions are grounded in text and tables that are extracted from the earnings reports of S&P 500 companies. Table 5 shows a QA data sample in FinQA.
The original FinQA dataset was constructed by financial professionals who knew the full document context for each question, but the released data only contains select content from a single page of the document. We extend this by providing the full SEC filing for each question, resulting in a more challenging task for modern LLMs. Additionally, the original task was to provide a numerical answer for each question. This provides no transparency for how the model arrived at its answer, which can limit user trust Ehsan et al. (2019); Lamm et al. (2020). We follow the setup of Koncel-Kedziorski et al. (2023) by requiring the model to generate executable Python code to produce an answer. This provides insight into how a model produced its answer – specifically, via illustrating what information from the context is used, along with what arithmetic operations are performed.
Dataset Representation: Each question in FinQA is a triplet $(c^{golden},q,a)$ of a golden context $c^{golden}$, a question $q$, and an answer $a$ written in human language. We extend the dataset in two ways: (1) context $c^{golden}$ is extended to the full document context $D$; and (2) we added a Python program $p$ that produces the answer $a$. Each final sample in DocFinQA is a quartet $(D,q,p,a)$. An example of this can be seen in Appendix J.
Filings Collection: For each question of the FinQA dataset, we identify the corresponding SEC filing from which it was originally annotated. We retrieve the full filing in HTML/XML format from SEC EDGAR service [^1] https://www.sec.gov/edgar/. Figure 2 shows the distribution of document length in DocFinQA. Since each filing contains many tables, maintaining the structure and order during extraction is critical for the numerical reasoning task. We convert each HTML-formatted filing to PDF format and use Kensho Extract [^2] https://kensho.com/extract to parse the filing into text format with tables being represented in markdown format. This process ensures that: (i) our dataset is rooted in authentic and relevant financial documentation and (ii) all the tables in the filings are parsed with high precision into a consistent format without any HTML-tag noise. We discuss our parsing choices in more detail in Appendix D.
Figure 2: Histogram of document length (#words) in DocFinQA dataset. Dash line shows the average length.
Chunking and Alignment: To study retrieval-based QA systems, we split each document $D$ into a set of chunks $C=\{c_{1},\cdots,c_{n}\}$. Each chunk contains 2,750 characters (equivalent to approximately 509 tokens) with 20% overlapping to avoid missing important context at the edges. To compute the performance, from the chunk set $C$ associated with each document $D$, we identify the best context chunk, $c^{*}$ that includes the information to answer question $q$. Since FinQA already provides the $c^{golden}$, we compute pair-wise scores between the chunks in $C$ and the golden chunk $c^{golden}$. We found that four-gram based similarity score offers the sharpest matching signal among tri-gram, four-gram, and fuzzy matching. The chunk that yields the highest score is selected as the target context chunk for retrieval.
We verify that this process results in good $c^{*}$ chunks through manual inspection and by substituting $c^{*}$ for $c^{golden}$ in a few-shot QA evaluation. This evaluation using GPT-3.5 on the FinQA test set shows highly similar performance (67.5% vs 67.3% accuracy).
Code Generation: The FinQA dataset provides solutions in a “program” syntax that, when executed, yields the golden result (e.g., subtract(34.8, 1.2), divide(#0, 34.8)). However, this derivation can be difficult to interpret and provides no meaningful context of what was being calculated. Koncel-Kedziorski et al. (2023) augments FinQA with readable Python code that can be executed to derive the answer, providing a layer of reliability and interpretability for the task at hand. Thus, we use the code-enhanced version of FinQA to construct DocFinQA (See Appendix F).
Statistics: The resultant DocFinQA dataset comprises 5,798 training, 791 development, and 1,032 test samples, derived from a diverse set of 1,236 unique SEC filings. By encapsulating a broad temporal scope, our dataset not only addresses the immediate need for long-context financial data but also aligns with the evolving dynamics of the financial landscape. Table 1 shows the statistics and characteristics of DocFinQA in comparison with other finance numerical reasoning and long-document QA datasets.
Due to the limited capability of the LLM we used to generate the code and the collection of the SEC filings, DocFinQA use a subset of FinQA questions. We analyze the distribution of the question types of FinQA and DocFinQA datasets (See Appendix A). We do not see any loss of any major question types in the dataset.
4 Retrieval-based QA Evaluation
Retrieval Task We test three encoding models for context retrieval: ColBERT (ColB.)Khattab and Zaharia (2020), Sentence-BERT (SentB.) Reimers and Gurevych (2019), and OpenAI’s Ada (Ada) Greene et al. (2022). Further, we finetune the ColBERT model (FT ColB.) on the training set of DocFinQA to measure performance using an in-domain model. We explored matching-based models (e.g., BM25 method Robertson et al. (1995)) but observe poor performance (See Appendix H).
To retrieve context for a question $q$ over chunk set $C=\{c\}$, we encode both $q$ and $C$ with an encoding model. This results in a set of embeddings for the each chunk, $V(c_{i})$ and an embedding for the question, $V(q)$. We compute the cosine similarity between $V(c_{i})$ and $V(q)$ to retrieve the top-k most similar chunks, $\{r_{1},r_{2},..r_{k}\}$.
Figure 3: Hit rate of retrieval models.
We evaluate these models using hit rate (HR@k). We compute the HR using target $c^{*}$ on the test set of DocFinQA. The results are shown in Figure 3. The Finetuned ColBERT model yields the highest HR, followed by the Original ColBERT model. The finetuned ColBERT model yields an average improvement of 91% HR over Sentence-BERT. We obtain a maximum 0.35 (HR@1) and 0.55 (HR@3).
Question Answering Task We formulate the QA task as a few-shot in-context learning task Brown et al. (2020). Due to the limited context length, the context of each demonstration contains a single golden chunk, while the context of the query question may contain up to $k$ chunks. A sample prompt template is given in Figure 10.
Table 2 reports the performance of 11 state-of-the art models with low context length. Larger models perform better than smaller models, as expected (e.g., MPT 30B vs MPT 7B). Models trained on code yield higher accuracy than non-code models (e.g., CodeLlama vs Llama). Models with additional supervised finetuning (e.g., LLama 2/7B+SFT) and instruction tuning (e.g., GPT-3.5) are the best among the examined models. Notably, Mistral 7B outperforms several larger models, although it lags behind Llama 2/7B+SFT and GPT-3.5.
Finetuned ColBERT is the best retrieval model in all but one setting. It yields a marginal but consistent improvement over the Original ColBERT, and a large improvement over Sentence BERT and OpenAI ADA.
Model/Size ColB. SentB. ADA FT ColB.
Falcon/7B   2.31   0.30   1.21   1.82
MPT/7B   4.63   2.71   3.83   4.77
MPT/30B 17.32 11.11 12.12 18.07
Llama 2/7B 13.49 8.86 11.11 13.50
Llama 2/13B 18.73 12.68 14.94 19.13
CodeLlama/7B 15.60 12.18 15.15 16.75
CodeLlama/13B 19.10 13.79 18.78 21.01
Mistral/7B 23.16 14.90 21.51 24.97
Llama 2/7B+SFT 32.93 24.77 34.34 36.14
GPT-3.5/- 41.59 33.78 36.43 42.64
Table 2: Performance of the LLMs on DocFinQA test set. For each row, the best performance among all retrieval models is in bold. The fewshot setting is selected based on the best performance on the development set (See Appendix I).
5 Retrieval-free QA Evaluation
Recently developed LLMs can handle context lengths over 100k tokens. We explore the promise of these models for retrieval-free long document QA using the DocFinQA dataset.
Since the average size of documents in our data is still longer than the context window, we adopt different strategies for answering questions over short (fewer than 100K tokens) and longer documents. For short documents, we provide examples from FinQA dataset (i.e., with  700 words of context) as a few-shot prompt followed by the whole target document and question. For longer documents, we adapt the S2A strategy proposed in Weston and Sukhbaatar (2023), which summarizes relevant information from the document before answering the question. For each 100K chunk of the long document, we include question/context pairs from FinQA as examples of information to extract for a summary. We then input the generated summary to the model with the same prompt used with short documents. Due to budget constraints, we sub-sample 400 questions from the test set for conducting experiments with GPT4. We refer to this set as the “test subsample”. In our test set, we maintain an equal number of short and longer documents (i.e., 200 datapoints of each type).
Model w/ Retrieval Accuracy
GPT-4 No 34.8
    Short-only No 46.5
    Long-only w/ S2A No 23.0
    Long-only w/ Interactive No 20.0
Mistral/7B Yes 28.4
LLama 2/7B+SFT Yes 28.6
GPT-3.5 Yes 42.4
GPT-4 Yes 51.5
Table 3: Performance of LLMs on DocFinQA test set.
Table 3 shows a comparison of GPT-4 performance on the test subsample when given full document context. The performance on short documents is twice as good as over longer documents, but is still less than 50% accurate. This indicates that there is much room for improvement with document-level financial QA.
The difference between long and short document accuracy also implies that better methods are needed for combining information across multiple calls to a document-processing LM. In this work, we explored two options: the S2A setup described above, and an Iterative refinement approach where the output program is written and refined as the LLM processes each 100k section of the document. Here, a temporary answer program (initially “None”) is passed along with the input each time the LLM is called. This method proved slightly less performant than S2A, achieving 20% accuracy compared to 23% on the longer-document subset, respectively.
6 Conclusion
This paper introduces a highly practical setting for grounded document-level question answering over financial reports. Previous work provides select content, whereas we incorporate the full financial report — averaging 123k words — for each financial question. Our findings reveal that this more realistic setting presents a significantly more difficult challenge, thereby opening new avenues for research in quantitative financial question answering.
Appendix A Impact of Data Selection
DocFinQA was subsampled from the full FinQA dataset based on the correctness of the code produced by an open-source LLM. This process may potentially filter out a collection of question types that were not answered by the LLM due to its limited capability. This section investigates the impact of this process by comparing the distribution of the question types in FinQA and DocFinQA. To do this, we show the distribution of questions grouped by their first 2 non-stop words in the Figure 4. The most important observation is that, overall, the distribution of the question set in DocFinQA and FinQA are very similar. There are no major groups being filtered out by our data selection process. The dominant questions (above 1% in FinQA) remains dominant and no major impact on the percentages of those questions is observed. The mid group (above 0.2% in FinQA) question sets see a mixed effect. A large portion of these questions see an increase in percentage while some experience significant loss (e.g., “what percentual” and “what decrease”). Lastly, the long tail group ( under 0.2% in FinQA) either remains the same (e.g., “percent total” and “what greatest”) or is completely wiped out due to small population (e.g., “was average”, and “what return”).
Figure 4: Distribution of question grouped by question types in the original FinQA and DocFinQA. The x-axis (percentage) is presented in log-scale to magnify the differences between the two sets.
Appendix B Golden Chunk Position
Figure 5 shows the distribution of the position golden chunk with the documents. We see that most of the golden chunks appear within the first 250 chunks (approximately 125K tokens which can be fed into the newest generative models). Nonetheless, there are a substantial number of questions that the golden chunk appear beyond this threshold.
Figure 5: Histogram of position of the FinQA context in the original SEC filing that is split into chunks of size 2750.
Appendix C Model Details
In this work, we used the base models of Falcon, MPT, Llama 2, CodeLlama, and Mistral throughout our work. These models were not trained with supervised finetuning or reinforcement learning human feedback. The GPT-3.5 model employed in this study is gpt-3.5-turbo-0613 while the GPT-4 model used is gpt-4-1106-preview.
We also included the Llama 2/7B + SFT that was finetuned on the training set of DocFinQA with golden chunk from FinQA ($c^{golden}$). The finetuning process takes 3 epochs with a batch size of 32. We use the context provided by the FinQA dataset as the input due to the limited maximum token length of the model. The maximum token length is set to 2048. The model is finetuned on 8 x Nvidia A100-80GB GPUs. We use AdamW optimizer with learning rate of 2e-6.
Appendix D Parsing SEC Filings
We explore different methods for parsing SEC filings consisting of HTML and XML markup into text and markdown tables for use in our QA systems. To evaluate parsing strategies, we measure HR@k when searching for the gold chunk among all document chunks for a single document using the FinQA question as the search query. Queries and document chunks are encoded with OpenAI’s ADA model. We compare BeautifulSoup, a standard library for manipulating HTML and XML formatted data, and Kensho Extract, a finance specific text and table extraction model. [^3] Passing the raw HTML/XML to the language model produces near-zero performance.
Figure 6 shows the performance of these two methods. We see that the finance specific models used in Kensho Extract result in better downstream performance of this retrieval setup compared to Beautiful Soup. Qualitative analysis of the different parsers reveals that Kensho Extract is better at structuring the tables used in financial documents, resulting in better readability which seems to extend to the encodings.
Figure 6: Accuracy for varying HR@ for two context extraction methods.
Appendix E ColBERT Finetuning
We finetune the original ColBERT v1 model on the train set of DocFinQA. For each data point we perform chunking and alignment to generate one golden chunk and $n-1$ negative chunks. For training, we generate a list of tuples (qid, pid+, pid-), where qid refers to the question, pid+ refers to the golden chunk and pid- refers to each of the negative chunks in that document. We train the model for a total of 3 epochs and store the checkpoints at the end of each epoch. The hit rate of the Finetuned ColBERT model after each epoch on the development set is shown in Figure 7. We observe that after the first epoch, additional finetuning does not show any performance improvement. The Finetuned ColBERT model referred to in this study thus uses the weights after the first epoch of training.
Figure 7: Hit rate of different ColBERT variants on the development set of DocFinQA.
Appendix F Code Conversion
Figure 8 shows the steps of converting (a) derivation of the result in FinQA into (b) dummy Python code with dummy variable names , and finally transform it to (c) a meaningful Python program in DocFinQA following the work by Koncel-Kedziorski et al. (2023).
(a) $subtract(34.8,1.2),divide(\#0,34.8)$ (b) $a=34.8-1.2$ $b=a/34.8$ $c=b*100$ (c) $payments\_decrease=34.8-1.2$ $change=payments\_decrease/34.8$ $answer=change*100$
Figure 8: Example of code conversion. (a) Original FinQA’s derivation. (b) Dummy Python Program (c) Meaningful Python Code in DocFinQA.
Appendix G Few-shot Settings
Due to the limited context length of the LLMs, the number of few-shot demonstrations and the number of chunks fed into the In-Context Learning must be optimized. We explore 3 settings of number of few-shot examples and 4 settings of number of chunks used as context in the query. Figure 9 shows the performance of these settings in with retrieval and answered by LLama 13B and CodeLLaMa 13B on the development set. We see that a higher number of few-shot example (numshot=3) yield consistent better performance compared to a lower one (numshot=1).
Figure 9: A QA performance plot on the development set of DocFinQA for the Llama 2 13B and CodeLlama 2 13B models for each of the 12 configurations
Figure 10 shows a the prompt template with in-context learning that we used.
Context: {golden chunk} Question: {question} Python Program: {program} Answer: {answer} Context: {golden chunk} Question: {question} Python Program: {program} Answer: {answer} Context: {golden chunk} Question: {question} Python Program: {program} Answer: {answer} Context: {first chunk} {second chunk} {third chunk} Question: {question} Python Program:
Figure 10: Prompt template with Top-3 context and 3-shot In-Context Learning.
Appendix H Detailed performance of retrieval method
Figure 11 shows a pilot study comparing dense-retrieval with OpenAI ADA and Sentence BERT versus sparse retrieval (BM 25) on the development set. We can clearly see that dense retrieval model offer a much higher hit ratio.
Figure 11: Accuracy for varying HR@ for three search methods on the development set
Appendix I Detailed performance on dev set
Table 4 reports the full performance on the development set with four retrieval models and three few-shot settings. This results in a total of twelve unique configurations. For our analysis we employ ten state-of-the-art LLMs and a finetuned version of LLama2 7B. For both finetuned and pretrained models we use greedy decoding whenever applicable. For hyperparameter tuning, we carry out the extraction task on the development set of DocFinQA. One trend noted was that all generic LLMs showed higher accuracy with shorter context and more fewshot example i.e. top chunk with 3 shot. While the code based LLMs such as Starcoder and CodeLLama showed higher accuracy with longer context i.e top 3 chunks with 3 shot. This trend is also depicted in Figure 9.
Model Size Upper Bound Original ColBERT Finetuned ColBERT Sentence-BERT OpenAI ADA
* * Top 1 Top 3 Top 3 Top 1 Top 3 Top 3 Top 1 Top 3 Top 3 Top 1 Top 3 Top 3
1 shot 3 shot 3 shot 1 shot 3 shot 3 shot 1 shot 3 shot 3 shot 1 shot 3 shot 3 shot 1 shot 3 shot
Falcon 7B 2.0 2.0 1.9 0.0 0.0 1.9 1.3 0.0 1.2 0.1 1.3 2.0 0.1 0.0
MPT 7B 6.8 6.6 4.5 0.8 0.2 4.9 1.0 1.2 3.9 0.6 0.8 4.3 1.6 2.0
MPT 30B 27.1 31.0 15.3 2.2 1.7 16.8 3.2 3.8 1.1 3.8 2.7 15.7 10.4 5.1
Llama 2 7B 17.3 22.0 12.8 5.8 8.0 14.0 6.0 10.3 8.9 2.7 6.5 11.2 4.0 11.0
Llama 2 + SFT 7B 67.1 69.7 30.0 32.6 31.3 32.2 35.3 33.9 19.9 24.1 24.3 28.7 29.4 27.7
Llama 2 13B 30.0 33.4 14.4 10.4 14.1 19.1 11.9 14.5 14.9 7.9 10.2 18.3 9.8 13.7
CodeLlama 7B 26.9 34.0 12.6 11.4 16.1 15.7 12.3 16.8 11.9 8.9 13.2 15.4 14.2 17.5
CodeLlama 13B 32.1 39.0 19.5 14.8 21.5 21.2 15.7 22.5 13.2 8.5 16.0 18.3 14.4 20.9
Mistral 7B 39.7 48.8 23.0 18.8 21.3 25.9 16.8 25.2 19.0 13.6 17.6 20.9 18.8 22.1
GPT 3.5 - 67.3 67.5 36.0 39.0 38.8 38.8 40.7 40.2 24.8 30.1 36.3 35.0 36.5 36.9
Table 4: Performance of the models on DocFinQA in one-shot and few-shot in-context learning settings for top 1 and top 3 retrieved chunk contexts on the development set. For each model, the best performance among all configurations is in bold. For each model, the best performance among different configurations for the same retrieval model is highlighted in violet. Top 1 and Top 3 indicate the number of retrieved chunks used as context for a configuration. *The single original context chunk from FinQA test set is used to estimate the upper bound.
Appendix J Prompt examples
Table 5 and J shows examples of context, question, and model output in FinQA and DocFinQA.
Context:
Performance graph comparison of five-year cumulative total return the following graph and table compare the cumulative total return on Citi 2019s common stock, which is listed on the NYSE under the ticker symbol 201cc 201d and held by 65691 common stockholders of record as of January 31, 2018, with the cumulative total return of the S&P 500 index and the S&P financial index over the five-year period through December 31, 2017. The graph and table assume that \$ 100 was invested on December 31, 2012 in Citi 2019s common stock, the S&P 500 index and the S&P financial index, and that all dividends were reinvested . comparison of five-year cumulative total return for the years ended date Citi S&P 500 financials.
| DATE | CITI | S&P 500 | S&P FINANCIALS |
| 31-Dec-2012 | 100.0 | 100.0 | 100.0 |
| 31-Dec-2013 | 131.8 | 132.4 | 135.6 |
| 31-Dec-2014 | 137.0 | 150.5 | 156.2 |
| 31-Dec-2015 | 131.4 | 152.6 | 153.9 |
| 31-Dec-2016 | 152.3 | 170.8 | 188.9 |
| 31-Dec-2017 | 193.5 | 208.1 | 230.9 |
Question:
What was the percentage cumulative total return for the five year period ended 31-dec-2017 of citi common stock?
Answer:
93.5%
Table 5: Example from FinQA dataset. The context provided here has been formatted from on the original dataset values
Context:
UNITED STATES SECURITIES AND EXCHANGE COMMISSION
# ANNUAL REPORT PURSUANT TO SECTION 13 OR 15(d) OF THE SECURITIES EXCHANGE ACT of 1934
For the Fiscal Year Ended December 30, 2006
Commission file number 1-4171
# Kellogg Company
(Exact Name of Registrant as Specified in its Charter)
Delaware (State of Incorporation) (I.R.S. Employer Identification No.) One Kellogg Square (Address of Principal Executive Offices) Securities registered pursuant to Section 12(b) of the Securities Act: Title of each class: Name of each exchange on which registered:
$\cdots$
The Consolidated Financial Statements and related Notes, together with Management’s Report on Internal Control over Financial Reporting, and the Report thereon of Pricewaterhouse Coopers LLP dated February 23, 2007, are included herein in Part II, Item 8.
# (a) 1. Consolidated Financial Statements Consolidated Statement of Earnings for the years ended December 30, 2006, December 31, 2005 and January 1, 2005. Consolidated Statement of Shareholders’ Equity for the years ended December 30, 2006, December 31, 2005 and January 1, 2005. Notes to Consolidated Financial Statements.
# (a) 2. Consolidated Financial Statement Schedule All financial statement schedules are omitted because they are not applicable or the required information is shown in the financial statements or the notes thereto.
# (a) 3. Exhibits required to be filed by Item 601 of Regulation S-K The information called for by this Item is incorporated herein by reference from the Exhibit Index on pages 61 through 64 of this Report. Pursuant to the requirements of Section 13 or 15(d) of the Securities Exchange Act of 1934, the Registrant has duly caused this Report to be signed on its behalf by the undersigned, thereunto duly authorized, this 23rd day of February, 2007. Pursuant to the requirements of the Securities Exchange Act of 1934, this Report has been signed below by the following persons on behalf of the Registrant and in the capacities and on the dates indicated. Electronic(E), | 10.48 | | IBRF |
| | Commission file number 1-4171.* | |
| 21.01 | Domestic and Foreign Subsidiaries of Kellogg. | E |
| 23.01 | Consent of Independent Registered Public Accounting Firm. | E |
| 24.01 | Powers of Attorney authorizing Gary H. Pilnick to execute our Annual Report on Form 10-K for the fiscal year ended December 30, 2006, on behalf of the Board of Directors, and each of them. | E |
| 31.1 | Rule 13a-14(a)/15d-14(a) Certification by A.D. David Mackay. | E |
| 31.2 | Rule 13a-14(a)/15d-14(a) Certification by John A. Bryant. | E |
| 32.1 | Section 1350 Certification by A.D. David Mackay. | E |
| 32.2 | Section 1350 Certification by John A. Bryant. | E |
Question:
What was the average cash flow from 2004 to 2006?
Program:
    net_cash_2006 = 957.4
    net_cash_2005 = 769.1
    net_cash_2004 = 950.4
    total_net_cash = net_cash_2006 + net_cash_2005 + net_cash_2004
    average_net_cash = total_net_cash / 3
    answer = average_net_cash
Answer:
892.3
Table 6: Examples from DocFinQA dataset with text and tables from entire SEC document as context (truncated for legibility), question, associated program and answer. A full report can be founded here https://www.annualreports.com/HostedData/AnnualReportArchive/k/NYSE_K_2006.pdf