### v2

# DocFinQA: A Long-Context Financial Reasoning Dataset

Abstract
For large language models (LLMs) to be effective in the financial domain – where each decision can have a significant impact – it is necessary to investigate realistic tasks and data. Financial professionals often interact with documents that are hundreds of pages long, but most financial research datasets only deal with short excerpts from these documents. To address this, we introduce a long-document financial QA task. We augment 7,437 questions from the existing FinQA dataset with the full-document context, extending the average context length from under 700 words in FinQA to 123k words in DocFinQA. We conduct extensive experiments over retrieval-based QA pipelines and long-context language models. DocFinQA proves a significant challenge for even state-of-the-art systems. We also provide a case-study on the longest documents in DocFinQA and find that models particularly struggle on these documents. Addressing these challenges may have a wide reaching impact across applications where specificity and long-range contexts are critical, like gene sequences and legal document contract analysis. The data and code is publicly accessible at github.com/anonymous.
1 Introduction
The frequent need to reason over large volumes of textual and tabular data makes financial analysis particularly challenging for LLMs Azzi et al. (2019). Existing work on automating financial numerical reasoning focuses on unrealistically specific document snippets Chen et al. (2021); Zhu et al. (2021). Datasets are often limited to pre-selected document sections, failing to reflect the broader and more realistic scenarios faced by analysts Masson and Montariol (2020). Financial professionals often sift through hundreds of pages per document, requiring deep understanding of both the content and structure to effectively navigate and extract pertinent information. Current long-document QA datasets such as NarrativeQA Kočiský et al. (2018) do not test the quantitative reasoning skills needed in the financial domain.
In this work, we introduce DocFinQA, a long-document financial question answering task. We extend the FinQA dataset of expert annotated questions and answers Chen et al. (2021) with full Securities and Exchange Commission (SEC) reports. This results in a significantly longer context in the DocFinQA dataset – by a factor of 175 – than the FinQA dataset. The resulting long-document QA task offers a more realistic evaluation of a model’s reasoning capabilities over financial documents. In line with recent work on program synthesis for financial QA Koncel-Kedziorski et al. (2023), the questions in DocFinQA are annotated with Python programs to generate the answers, allowing for training and evaluating program synthesis models for use in realistic financial workflows.
Figure 1: DocFinQA extends FinQA to documents often over 150 pages long (100K+ tokens), so it is difficult to find the pertinent information. The question for the example above is: “For the quarter December 31, 2012 what was the percent of the total number of shares purchased in December?” The correct answer is 16.5%.
Using this setup, we evaluate retrieval-based and long-context LLM systems. We study a typical retrieval pipeline that chunks and encodes the document, searching for the best chunks given a question, and passing the question and top-$k$ chunks to a generative QA model Hsu et al. (2021).
We also evaluate retrieval-free approaches using long-context LLMs Weston and Sukhbaatar (2023). Our results show that the successful employment of LLMs in financial settings requires further study of the specific nuances of the financial domain, such as context disambiguation. Our dataset represents a step towards better capturing these nuances.
2 Related Work
Prior studies in financial question answering focus on non-numerical reasoning Day and Lee (2016); Jørgensen et al. (2023); Maia et al. (2018). Short-context grounded numerical reasoning tasks were introduced with datasets such as FinQA Chen et al. (2021) and TAT-QA Zhu et al. (2021). Recently, understanding long documents has attracted more attention for tasks involving events Yang et al. (2018), table of contents Bentabet et al. (2020), and causal relations Mariko et al. (2022). However, to the best of our knowledge, this is the first attempt to address financial numerical QA by grounded in long documents with upwards of hundreds of pages of context for each question.
Long-document QA has been studied in NLP with the introduction of datasets such as SearchQA Dunn et al. (2017), NarrativeQA Kočiský et al. (2018), QuALITY Pang et al. (2022), and PDFTriage Saad-Falcon et al. (2023). Due to the limited context size of LLMs, retrieval-based models are commonly used to filter irrelevant text Izacard et al. (2022); Lewis et al. (2020). Recently, advances in attention mechanisms Beltagy et al. (2020); Dao et al. (2022) and positional embeddings Press et al. (2021); Su et al. (2023) allow for end-to-end grounded QA with context windows of more than 100k tokens. However, these methods suffer from loss of important context Zhang et al. (2023) and often fail to make full use of longer inputs Liu et al. (2023). Our work studies the intersection of numerical reasoning and long-document processing, and our results demonstrate that there is still ample room for improvement in this domain.
Dataset #Docs #QAs #Words Pag Num Tab
[…]
Table 1: Comparison of DocFinQA and existing Finance QA and Long Document QA dataset. DocFinQA includes multi-page documents rich with both numeric data and tables.
3 DocFinQA Dataset
Our dataset is an extension of the FinQA dataset. (See Table 5 for an example of FinQA). FinQA was created by finance experts from S&P 500 companies’ annual financial report (aka., 10-K). They used full documents when creating the questions (e.g., 100s of pages), but FinQA only includes the curated context necessary to produce correct answers (e.g., a paragraph and/or table of statistics).
To recreate the original problem, we retrieve the full SEC filing for each question. Following Koncel-Kedziorski et al. (2023), we require the model to generate Python code to produce an answer instead of directly generating the numeric answer. The generated Python code illustrates what information from the context is used, along with what arithmetic operations are performed.
Dataset Representation: Each question in FinQA is a triplet $(c^{golden},q,a)$ composed of a golden context $c^{golden}$, a question $q$, and an answer $a$ written in human language. We extend the dataset in two ways: (1) context $c^{golden}$ is extended to the full document context $D$; and (2) we added a Python program $p$ that produces the answer $a$. Each final sample in DocFinQA is a quartet $(D,q,p,a)$. See Appendix M for an example.
Filings Collection: For each question of the FinQA dataset, we identify the corresponding SEC filing from which it was originally created. We retrieved the filing in HTML/XML format from SEC’s EDGAR service and parsed text and table into clean markdown format text Wang et al. (2023). The collection and parsing processes are presented in more detail in Appendix F and Appendix G, respectively. Figure 2 shows the distribution of document lengths in DocFinQA.
Figure 2: Histogram of document length (#words) in DocFinQA dataset with dash line representing the average length of the documents. Purple line depicts the proportion of documents where the question context is within the current number of words.
Chunking and Alignment: To study retrieval-based QA systems, we split each document $D$ into a set of chunks $C=\{c_{1},\cdots,c_{n}\}$. Each chunk contains 2,750 characters ($\sim\!509$ tokens) with a 20% overlap to avoid missing context at the edges. To compute the performance, we identify the best context chunk, $c^{*}$, from the chunk set $C$ associated with each document $D$ that includes the information to answer question $q$. Since FinQA already provides $c^{golden}$, we compute a pair-wise score $(c_{i},c^{golden})$, for all chunks, $c_{i}\in C$, including the golden chunk. We found that four-gram based similarity score offers the sharpest matching signal among tri-gram, four-gram, and fuzzy matching. The chunk with the highest score is selected as the target context chunk for retrieval. We verify that this process results in good $c^{*}$ chunks through manual inspection and by substituting $c^{*}$ for $c^{golden}$ in a few-shot QA evaluation with GPT-3.5. The model scores slightly higher when using $c^{*}$ rather than $c^{golden}$ (67.5% vs 67.3% accuracy respectively), suggesting that little context is lost.
Code Generation: The FinQA dataset provides solutions in a “program” syntax that, when executed, yields the answer (e.g., in Figure 1 the solution is divide(102400, 619314). However, this derivation does not provide meaningful context of what is being calculated. In our running example, 102400 is not semantically grounded to the document. Koncel-Kedziorski et al. (2023) augments FinQA with readable Python code (including named variables like, dec_shares = 102_400) that can be executed to derive the answer, providing a layer of interpretability. Thus, we use the code-enhanced version of FinQA to construct DocFinQA (See Appendix I).
Statistics: The resultant DocFinQA dataset comprises 5,735 training, 780 development, and 922 test samples, derived from 801 unique SEC filings. Table 1 shows the statistics and characteristics of DocFinQA in comparison with other finance numerical reasoning and long-document QA datasets.
Due to the limited availability of complete SEC filings (refer Appendix F) and imperfections in the code generation process, DocFinQA encompasses a subset of FinQA questions. We analyze the distribution of the question types of FinQA and DocFinQA datasets (See Appendix C). DocFinQA remains a representation of all major question types from the original FinQA dataset.
4 Retrieval-based QA Evaluation
Retrieval Task We test three models for context retrieval: ColBERT (ColB)Khattab and Zaharia (2020), Sentence-BERT (SentB) Reimers and Gurevych (2019), and OpenAI’s Ada Greene et al. (2022). Further, we finetune the ColBERT model (FT ColB) on the training set of DocFinQA to evaluate an in-domain model. We test a matching-based model, BM25 Robertson et al. (1995), but observe poor performance (see Appendix K).
To retrieve context for a question $q$ over chunk set $C$, we encode both $q$ and $C$ with the encoding models mentioned above. This results in an embedding for the question $v_{q}$ and chunk embeddings $V_{C}=\{v_{c_{i}}|c_{i}\in C\}$. We compute the cosine similarity between $v_{q}$ and each vector in $V_{C}$ to retrieve the top-k most similar chunks. We evaluate these models using hit rate (HR@k) on the test set of DocFinQA using the target $c^{*}$. Results are shown in Figure 3. The FT ColB yields the highest HR, followed by ColB. FT ColB yields an average improvement of 91% HR over SentB and obtains a 0.35 (HR@1) and 0.55 (HR@3).
Question Answering Task We formulate the QA task as a few-shot in-context learning task Brown et al. (2020). For each in-context example, we only provide the gold chunk and the correct answer. For the actual query we provide $k$ chunks. A sample prompt template is given in Figure 10.
We evaluate Falcon Penedo et al. (2023), MPT MosaicML (2023), LLaMa 2 and CodeLlaMa Touvron et al. (2023), Mistral Jiang et al. (2023), GPT3.5 Brown et al. (2020); OpenAI (2023) models. We weren’t able to evaluate proprietary models such as GPT3 Brown et al. (2020), BloombergGPT Wu et al. (2023) due to their inaccessibility. We skipped models that were not finetuned for code generation such as PIXIU Xie et al. (2023) and FinGPT Yang et al. (2023) due to their poor performance. We also skipped models trained for other languages such as BBT-Fin Lu et al. (2023) and XuanYuan 2.0 Zhang and Yang (2023).
Table 2 reports the performance of 11 state-of-the art models. Larger models outperform smaller models (e.g., MPT 30B vs MPT 7B). Models trained on code yield higher accuracy than non-code models (e.g., CodeLlama vs Llama). Models with additional supervised finetuning (e.g., LLama 2/7B+SFT) and instruction tuning (e.g., GPT-3.5) are among the best examined. Notably, Mistral 7B outperforms several larger models, although it lags behind Llama 2/7B+SFT and GPT-3.5.
The FT ColB model is the best retrieval model in all but one setting. It yields a marginal but consistent improvement over the ColB, and a large improvement over SentB and Ada.
Model/Size ColB SentB ADA FT ColB
Falcon/7B   2.3   0.3   1.2   1.8
MPT/7B   4.6   2.7   3.8   4.8
MPT/30B 17.3 11.1 12.1 18.1
Llama 2/7B 13.5 8.9 11.1 13.5
Llama 2/13B 18.7 12.7 14.9 19.1
CodeLlama/7B 15.6 12.2 15.2 16.8
CodeLlama/13B 19.1 13.8 18.8 21.0
Mistral/7B 23.2 14.9 21.5 25.0
Llama 2/7B+SFT 32.9 24.8 34.3 36.1
GPT-3.5/- 41.6 33.8 36.4 42.6
Table 2: Performance on DocFinQA test set. For each row, the best performance among all retrieval models is in bold. The fewshot setting is selected based on the best performance on the dev set (See Appendix L).
5 Case Study w/ 100K+ Token Documents
Recent LLMs can handle context lengths of 128K tokens, but more than 40% of the documents in DocFinQA remain unanswerable even at this content length (see Figure 2). Here, we evaluate performance on a test subsample of 200111The test subsample is limited to 200 documents due to the monetary and temporal costs of human evaluation and GPT4. long documents, each of which has 100K or more tokens.
We explore two retrieval-free options. System 2 Attention (S2A) extracts relevant information from each 100K-token chunk of a document before answering the question using the combined extracted information as context  (Weston and Sukhbaatar, 2023). The Iterative method produces the output program iteratively as the LLM processes each 100k section of the document. A temporary answer program (initially “None”) is input with each section to the LLM. We also report the performance of the best retrieval-based model (Retrieval) based on the experiment in Section 4. The human evaluation we conducted on these 200 questions highlights the challenging nature of this dataset. [^2] Experienced but non-expert human participants. See Appendix A for details.
Model/Size + Method w/ Retrieval Test Subsample
Human No 41.0
Mistral/7B + Iterative No 11.5
Mistral/7B + S2A No 15.5
Mistral/7B + Retrieval Yes 20.0
GPT-4 + Iterative No 20.0
GPT-4 + S2A No 23.0
GPT-4 + Retrieval Yes 47.5
Table 3: Retrieval-free performance on a case-study of $100K+$ token documents.
Non-expert human performance on DocFinQA is lower than FinQA (41% versus 50.7%). This can be attributed to the difficulty of also finding the golden page, compared to the golden page being given in FinQA. Notably, the expert performance reported in FinQA is 91.2%.
Still, non-expert human performance is double that of retrieval-free GPT-4 on these long documents, and roughly triple that of retrieval-free Mistral. Secondly, Iterative performed worse than S2A for both GPT-4 and Mistral with a reduced accuracy of 3% and 4%, respectively. Third, with retrieval, both Mistral and GPT-4 outperform their retrieval-free counterparts, with the assisted GPT-4 now on par with the human cohort. Together, these results highlight that DocFinQA is an effective and difficult test for long-document QA, and that there is still room for significant improvement in this domain. For instance, further exploration into methods that combine information across multiple calls to a document-processing LLM is warranted.
6 Conclusion
This paper introduces a realistic document-level question answering dataset over financial reports. Each question includes a full financial report (averaging 123K words), a far greater challenge than previous work that hones in on pre-specified content. Our findings reveal that this more realistic setting presents a significantly more difficult challenge, thereby opening new avenues for research in quantitative financial question answering.
7 Limitation
This work introduced an extension of the existing FinQA dataset. Due to limited human resources, we only validate the test set while the training and the development set were not fully validated. As a result, we can not make any claim of bias and question quality in the not-yet-validated data points offered in this paper. Additionally, as discussed in section 2, the code provided in this work was generated by WizardCoder LLMs. Our assumption is that the code is correct if it produces correct or approximately close to the golden answer. This method may generate both false positive codes (the code that generate correct answer with incorrect rationales) and false negative codes (the correct code that fail the approximation test).
8 Broader Impact and Ethical Considerations
We do not foresee any considerable risks associated with our work given that it is an extension of a publicly available documents and dataset. To uphold transparency, the paper provides detailed documentation of the dataset creation process, including the sources of data and annotation details. Our dataset serves as a resource to underscore the need for more long context oriented benchmarks both within and outside the financial domain and does not intend to criticize any one or more LLMs.
The annotation in this work is done automatically, so no crowd-source or contract annotators were hired throughout the process. The human evaluation in this study were done by full-time paid coworkers known to the authors.
Appendix A Human Evaluation Setting
We recruited three data professionals with 4-5 years of experience working with financial documents, including but not limited to 10-K filings, to estimate human evaluation. The professionals were provided with the entire document in PDF format, maintaining the SEC’s original format for ease of reading. They were allowed to use the keyword-search feature of PDF reader applications and a simple calculator for basic arithmetic operations required for this task. On average, the professionals spent 25 minutes per question.
Appendix B Retrieval Performance
Figure 3: Hit rate of retrieval models.
Appendix C Impact of Data Selection
DocFinQA was subsampled from the full FinQA dataset based on the correctness of the code produced by an open-source LLM. This process may potentially filter out a collection of question types that were not answered by the LLM due to its limited capability. This section investigates the impact of this process by comparing the distribution of the question types in FinQA and DocFinQA. To do this, we show the distribution of questions grouped by their first 2 non-stop words in the Figure 4. The most important observation is that, overall, the distribution of the question set in DocFinQA and FinQA are very similar. There are no major groups being filtered out by our data selection process. The dominant questions (above 1% in FinQA) remains dominant and no major impact on the percentages of those questions is observed. The mid group (above 0.2% in FinQA) question sets see a mixed effect. A large portion of these questions see an increase in percentage while some experience significant loss (e.g., “what percentual” and “what decrease”). Lastly, the long tail group ( under 0.2% in FinQA) either remains the same (e.g., “percent total” and “what greatest”) or is completely wiped out due to small population (e.g., “was average”, and “what return”).
Figure 4: Distribution of question grouped by question types in the original FinQA and DocFinQA. The x-axis (percentage) is presented in log-scale to magnify the differences between the two sets.
Appendix D Golden Chunk Position
Figure 5 shows the distribution of the position golden chunk with the documents. We see that most of the golden chunks appear within the first 250 chunks (approximately 125K tokens which can be fed into the newest generative models). Nonetheless, there are a substantial number of questions that the golden chunk appear beyond this threshold.
Figure 5: Histogram of position of the FinQA context in the original SEC filing that is split into chunks of size 2750.
Appendix E Model Details
In this work, we used the base models of Falcon, MPT, Llama 2, CodeLlama, and Mistral throughout our work. These models were not trained with supervised finetuning or reinforcement learning human feedback. The GPT-3.5 model employed in this study is gpt-3.5-turbo-0613 while the GPT-4 model used is gpt-4-1106-preview.
We also included the Llama 2/7B + SFT that was finetuned on the training set of DocFinQA with golden chunk from FinQA ($c^{golden}$). The finetuning process takes 3 epochs with a batch size of 32. We use the context provided by the FinQA dataset as the input due to the limited maximum token length of the model. The maximum token length is set to 2048. The model is finetuned on 8 x Nvidia A100-80GB GPUs. We use AdamW optimizer with learning rate of 2e-6. The training process takes 4 hours to complete.
Appendix F SEC Filing Collection
Each data point in the FinQA dataset consists of a document identification field as shown in Table 5. This field is made up of 3 sections separated by a forward slash. The first is a string called company ticker symbol, the second refers to the year in which this document was filed and the third is the page number in the document where the answer can be found.
Downloading the right 10-K filing from SEC begins with identifying the company code from the company ticker symbol. For example, C/2017/page_328.pdf-1 in FinQA maps to the CITIGROUP INC with company code 831001. This mapping is obtained from the official file released by SEC which can be found here https://www.sec.gov/file/company-tickers. We automatically generate a URL using the company code obtained. From the SEC website, either filings are downloaded as TXT, HTML or XBRL using the generated URL. At this stage, approximately 6.5% (or 543) data points which is corresponds to approximately 9.4% (or 17) documents were dropped, either due to lack of mapping or nonavailability of older documents. Further, conversion of the downloaded files to PDF caused a loss of 117 data points (19 unique documents) due to formatting issues.
Appendix G Parsing SEC Filings
Since each filing contains many tables, maintaining the structure and order during extraction is critical for numerical reasoning. We convert each HTML-formatted filing to PDF format and use a finance-specific PDF extractor to parse the filing into markdown format. This process ensures that: (i) our dataset is grounded in the relevant financial documentation and (ii) all the tables in the filings are parsed with high precision into a consistent format without any HTML-tag noise.
We explore different methods for parsing SEC filings consisting of HTML and XML markup into text and markdown tables for use in our QA systems. To evaluate parsing strategies, we measure HR@k when searching for the gold chunk among all document chunks for a single document using the FinQA question as the search query. Queries and document chunks are encoded with OpenAI’s ADA model. We compare BeautifulSoup, a standard library for manipulating HTML and XML formatted data, and Kensho Extract, a finance specific text and table extraction model. [^3] Passing the raw HTML/XML to the language model produces near-zero performance.
Figure 6 shows the performance of these two methods. We see that the finance specific models used in Kensho Extract result in better downstream performance of this retrieval setup compared to Beautiful Soup. Qualitative analysis of the different parsers reveals that Kensho Extract is better at structuring the tables used in financial documents, resulting in better readability which seems to extend to the encodings.
Figure 6: Accuracy for varying HR@ for two context extraction methods.
Appendix H ColBERT Finetuning
We finetune the original ColBERT v1 model on the train set of DocFinQA. For each data point we perform chunking and alignment to generate one golden chunk and $n-1$ negative chunks. For training, we generate a list of tuples (qid, pid+, pid-), where qid refers to the question, pid+ refers to the golden chunk and pid- refers to each of the negative chunks in that document. We train the model for a total of 3 epochs and store the checkpoints at the end of each epoch. The hit rate of the Finetuned ColBERT model after each epoch on the development set is shown in Figure 7. We observe that after the first epoch, additional finetuning does not show any performance improvement. The Finetuned ColBERT model referred to in this study thus uses the weights after the first epoch of training.
Figure 7: Hit rate of different ColBERT variants on the development set of DocFinQA.
Appendix I Code Conversion
Figure 8 shows the steps of converting (a) derivation of the result in FinQA into (b) dummy Python code with dummy variable names , and finally transform it to (c) a meaningful Python program in DocFinQA following the work by Koncel-Kedziorski et al. (2023).
(a) $subtract(34.8,1.2),divide(\#0,34.8)$ (b) $a=34.8-1.2$ $b=a/34.8$ $c=b*100$ (c) $payments\_decrease=34.8-1.2$ $change=payments\_decrease/34.8$ $answer=change*100$
Figure 8: Example of code conversion. (a) Original FinQA’s derivation. (b) Dummy Python Program (c) Meaningful Python Code in DocFinQA.
Appendix J Few-shot Settings
Due to the limited context length of the LLMs, the number of few-shot demonstrations and the number of chunks fed into the In-Context Learning must be optimized. We explore 3 settings of number of few-shot examples and 4 settings of number of chunks used as context in the query. Figure 9 shows the performance of these settings in with retrieval and answered by LLama 13B and CodeLLaMa 13B on the development set. We see that a higher number of few-shot example (numshot=3) yield consistent better performance compared to a lower one (numshot=1).
Figure 9: A QA performance plot on the development set of DocFinQA for the Llama 2 13B and CodeLlama 2 13B models for each of the 12 configurations
Figure 10 shows a the prompt template with in-context learning that we used.
Context: {golden chunk} Question: {question} Python Program: {program} Answer: {answer} Context: {golden chunk} Question: {question} Python Program: {program} Answer: {answer} Context: {golden chunk} Question: {question} Python Program: {program} Answer: {answer} Context: {first chunk} {second chunk} {third chunk} Question: {question} Python Program:
Figure 10: Prompt template with Top-3 context and 3-shot In-Context Learning.
Appendix K Detailed performance of retrieval method
Figure 11 shows a pilot study comparing dense-retrieval with OpenAI ADA and Sentence BERT versus sparse retrieval (BM 25) on the development set. We can clearly see that dense retrieval model offer a much higher hit ratio.
Figure 11: Accuracy for varying HR@ for three search methods on the development set
Appendix L Detailed performance on dev set
Table 4 reports the full performance on the development set with four retrieval models and three few-shot settings. This results in a total of twelve unique configurations. For our analysis we employ ten state-of-the-art LLMs and a finetuned version of LLama2 7B. For both finetuned and pretrained models we use greedy decoding whenever applicable. For hyperparameter tuning, we carry out the extraction task on the development set of DocFinQA. One trend noted was that all generic LLMs showed higher accuracy with shorter context and more fewshot example i.e. top chunk with 3 shot. While the code based LLMs such as Starcoder and CodeLLama showed higher accuracy with longer context i.e top 3 chunks with 3 shot. This trend is also depicted in Figure 9.
Model Size Upper Bound Original ColBERT Finetuned ColBERT Sentence-BERT OpenAI ADA
* * Top 1 Top 3 Top 3 Top 1 Top 3 Top 3 Top 1 Top 3 Top 3 Top 1 Top 3 Top 3
1 shot 3 shot 3 shot 1 shot 3 shot 3 shot 1 shot 3 shot 3 shot 1 shot 3 shot 3 shot 1 shot 3 shot
Falcon 7B 2.0 2.0 1.9 0.0 0.0 1.9 1.3 0.0 1.2 0.1 1.3 2.0 0.1 0.0
MPT 7B 6.8 6.6 4.5 0.8 0.2 4.9 1.0 1.2 3.9 0.6 0.8 4.3 1.6 2.0
MPT 30B 27.1 31.0 15.3 2.2 1.7 16.8 3.2 3.8 1.1 3.8 2.7 15.7 10.4 5.1
Llama 2 7B 17.3 22.0 12.8 5.8 8.0 14.0 6.0 10.3 8.9 2.7 6.5 11.2 4.0 11.0
Llama 2 + SFT 7B 67.1 69.7 30.0 32.6 31.3 32.2 35.3 33.9 19.9 24.1 24.3 28.7 29.4 27.7
Llama 2 13B 30.0 33.4 14.4 10.4 14.1 19.1 11.9 14.5 14.9 7.9 10.2 18.3 9.8 13.7
CodeLlama 7B 26.9 34.0 12.6 11.4 16.1 15.7 12.3 16.8 11.9 8.9 13.2 15.4 14.2 17.5
CodeLlama 13B 32.1 39.0 19.5 14.8 21.5 21.2 15.7 22.5 13.2 8.5 16.0 18.3 14.4 20.9
Mistral 7B 39.7 48.8 23.0 18.8 21.3 25.9 16.8 25.2 19.0 13.6 17.6 20.9 18.8 22.1
GPT 3.5 - 67.3 67.5 36.0 39.0 38.8 38.8 40.7 40.2 24.8 30.1 36.3 35.0 36.5 36.9
Table 4: Performance of the models on DocFinQA in one-shot and few-shot in-context learning settings for top 1 and top 3 retrieved chunk contexts on the development set. For each model, the best performance among all configurations is in bold. For each model, the best performance among different configurations for the same retrieval model is highlighted in violet. Top 1 and Top 3 indicate the number of retrieved chunks used as context for a configuration. *The single original context chunk from FinQA test set is used to estimate the upper bound.
Appendix M Prompt examples
Table 5 and M shows examples of context, question, and model output in FinQA and DocFinQA.
ID: C/2017/page_328.pdf-1
Context:
Performance graph comparison of five-year cumulative total return the following graph and table compare the cumulative total return on Citi 2019s common stock, which is listed on the NYSE under the ticker symbol 201cc 201d and held by 65691 common stockholders of record as of January 31, 2018, with the cumulative total return of the S&P 500 index and the S&P financial index over the five-year period through December 31, 2017. The graph and table assume that \$ 100 was invested on December 31, 2012 in Citi 2019s common stock, the S&P 500 index and the S&P financial index, and that all dividends were reinvested . comparison of five-year cumulative total return for the years ended date Citi S&P 500 financials.
| DATE | CITI | S&P 500 | S&P FINANCIALS |
| 31-Dec-2012 | 100.0 | 100.0 | 100.0 |
| 31-Dec-2013 | 131.8 | 132.4 | 135.6 |
| 31-Dec-2014 | 137.0 | 150.5 | 156.2 |
| 31-Dec-2015 | 131.4 | 152.6 | 153.9 |
| 31-Dec-2016 | 152.3 | 170.8 | 188.9 |
| 31-Dec-2017 | 193.5 | 208.1 | 230.9 |
Question:
What was the percentage cumulative total return for the five year period ended 31-dec-2017 of citi common stock?
Answer:
93.5%
Table 5: Example from FinQA dataset. The context provided here has been formatted from the original dataset values.
Context:
UNITED STATES SECURITIES AND EXCHANGE COMMISSION
# ANNUAL REPORT PURSUANT TO SECTION 13 OR 15(d) OF THE SECURITIES EXCHANGE ACT of 1934
For the Fiscal Year Ended December 30, 2006
Commission file number 1-4171
# Kellogg Company
(Exact Name of Registrant as Specified in its Charter)
Delaware (State of Incorporation) (I.R.S. Employer Identification No.) One Kellogg Square (Address of Principal Executive Offices) Securities registered pursuant to Section 12(b) of the Securities Act: Title of each class: Name of each exchange on which registered:
$\cdots$
The Consolidated Financial Statements and related Notes, together with Management’s Report on Internal Control over Financial Reporting, and the Report thereon of Pricewaterhouse Coopers LLP dated February 23, 2007, are included herein in Part II, Item 8.
# (a) 1. Consolidated Financial Statements Consolidated Statement of Earnings for the years ended December 30, 2006, December 31, 2005 and January 1, 2005. Consolidated Statement of Shareholders’ Equity for the years ended December 30, 2006, December 31, 2005 and January 1, 2005. Notes to Consolidated Financial Statements.
# (a) 2. Consolidated Financial Statement Schedule All financial statement schedules are omitted because they are not applicable or the required information is shown in the financial statements or the notes thereto.
# (a) 3. Exhibits required to be filed by Item 601 of Regulation S-K The information called for by this Item is incorporated herein by reference from the Exhibit Index on pages 61 through 64 of this Report. Pursuant to the requirements of Section 13 or 15(d) of the Securities Exchange Act of 1934, the Registrant has duly caused this Report to be signed on its behalf by the undersigned, thereunto duly authorized, this 23rd day of February, 2007. Pursuant to the requirements of the Securities Exchange Act of 1934, this Report has been signed below by the following persons on behalf of the Registrant and in the capacities and on the dates indicated. Electronic(E), | 10.48 | | IBRF |
| | Commission file number 1-4171.* | |
| 21.01 | Domestic and Foreign Subsidiaries of Kellogg. | E |
| 23.01 | Consent of Independent Registered Public Accounting Firm. | E |
| 24.01 | Powers of Attorney authorizing Gary H. Pilnick to execute our Annual Report on Form 10-K for the fiscal year ended December 30, 2006, on behalf of the Board of Directors, and each of them. | E |
| 31.1 | Rule 13a-14(a)/15d-14(a) Certification by A.D. David Mackay. | E |
| 31.2 | Rule 13a-14(a)/15d-14(a) Certification by John A. Bryant. | E |
| 32.1 | Section 1350 Certification by A.D. David Mackay. | E |
| 32.2 | Section 1350 Certification by John A. Bryant. | E |
Question:
What was the average cash flow from 2004 to 2006?
Program:
    net_cash_2006 = 957.4
    net_cash_2005 = 769.1
    net_cash_2004 = 950.4
    total_net_cash = net_cash_2006 + net_cash_2005 + net_cash_2004
    average_net_cash = total_net_cash / 3
    answer = average_net_cash
Answer:
892.3
Table 6: Examples from DocFinQA dataset with text and tables from entire SEC document as context (truncated for legibility), question, associated program and answer. A full report can be founded here https://www.annualreports.com/HostedData/AnnualReportArchive/k/NYSE_K_2006.pdf